Given an optimal capital structure that is 50% debt and 50% common stock, calculate the weighted average cost of capital for the company given the following additional information:

Bond coupon rate 8 %
Bond yield to maturity 5 %
Dividend, expected $ 5
Price, common $ 80
Growth rate 5 %
Corporate tax rate 21 %

Multiple Choice

  • More than 6% and less than 7%.

  • More than 8%.

  • More than 7% and less than 8%.

  • Less than 6%.
Answer #1

Cost of equity =Expected dividend/Current price+growth =5/80+5% =11.25%

Cost of debt =5%*(1-21%) =3.95%

WACC =(11.25%*50%)+(3.95%*50%) =7.6%

The answer is:

  • More than 7% and less than 8%.